OpenStudy (anonymous):

When Maggie applies the brakes to her car, it slows uniformly from 15.0 m/s to 0.0 m/s in 2.50 s. How many meters before a stop sign must she apply her brakes in order to stop at the sign?

OpenStudy (anonymous):

a=-15/2.5 v^2=u^2+2as 0=225-(300/25)s 12s=225 s=225/12